OpenStudy (anonymous):

You are given 100 Ω resistors. A. If in a series Rtotal= B. If in parallel Rtotal=

OpenStudy (anonymous):

two resistors?

OpenStudy (jtvatsim):

\[R_{total_{series}} = R_1 + R_2 + R_3 + \cdots\] \[\frac{1}{R_{total_{para}}} = \frac{1}{R_1} + \frac{1}{R_2} + \frac{1}{R_3} + \cdots\]

OpenStudy (anonymous):

in series, resistances add: \(R_{total}=\sum_i R_i\) in parallel, *conductances* (or inverse resistances) add: \(G_{total}=\sum_i G_i\) where \(G=1/R\) so \(\dfrac1{R_{total}}=\sum_i\dfrac1{R_i}\)

OpenStudy (anonymous):

So um what do you to know it? I only know it by Rtotalseries=R1+R2+R3+⋯ 1Rtotalpara=1R1+1R2+1R3+⋯

OpenStudy (jtvatsim):

Since we don't know how many resistors we have, let's just say we have N of them (that covers the bases). Each one is at 100, so Rtotalseries = 100 + 100 + 100 (N times) or = N(100) 1/Rtotalpara = 1/100 + 1/100 + 1/100 (N times) or = N(1/100) = N/100 so, Rtotalpara = 100/N.

OpenStudy (anonymous):

Oh so if it was 4 resistors then what?

OpenStudy (anonymous):

they gave you the formulas

OpenStudy (anonymous):



OpenStudy (anonymous):